# Semantic Structure-Aware Generative Attacks for Enhanced Adversarial Transferability

### v3

@@ 1 Introduction
Deep neural networks have driven transformative advances across computer vision, natural language processing, and medical diagnosis. Their capacity to learn rich, hierarchical feature representations has enabled breakthroughs in tasks ranging from fine-grained image classification to complex sequence generation. Yet these models remain alarmingly fragile: even small, human-imperceptible perturbations—known as adversarial examples (AEs) [58]—can provoke confident misclassification, underscoring the need to safeguard real-world deployments against malicious exploits. The threat is further intensified in black-box scenarios, where attackers have no access to a model’s architecture or parameters. In such scenarios, transferable adversarial attacks craft perturbations on a surrogate network and deploy the same perturbation against unknown targets. This capability empowers adversaries to compromise multiple real-world systems, such as autonomous vehicles, biometric scanners, and diagnostic tools, with a single perturbation strategy, posing severe risks to safety-critical applications.
@@
Early white-box iterative attacks, such as FGSM [85] and its iterative variants [15, 76, 17], exploit gradient information directly but inherently require full model access. To evaluate real-world attack robustness, transfer-based (black-box) attacks emerged, leveraging the phenomenon that adversarial perturbations crafted on one model often fool others. While iterative gradient-based methods [43, 9, 85], whether operating in white-box or query-based modes, can yield moderately transferable adversarial examples, they incur intensive costs due to per-sample optimization steps, resulting in prohibitive computation time and slow inference. Such bottleneck prohibits their use in time-sensitive scenarios and limits scalable and practically deployable attacks.
@@
To overcome this, generative transfer attacks [74, 68, 2, 3, 51, 50, 29, 86, 1, 81, 82, 83, 48] aim to train a dedicated perturbation generator against a surrogate, enabling fast inference and high transfer rates without per-example optimization. Generative attacks address these limitations by training a feed-forward generator to produce perturbations in one pass. Once trained, this generator yields adversarial noise with a single forward inference, offering orders-of-magnitude improvements in speed, seamless scalability to large datasets, and strong transferability. However, existing generative frameworks predominantly treat the generator as a black box, optimizing only end-to-end metrics while overlooking how semantic information, such as object edges and coarse shapes, is represented and potentially degraded within its intermediate layers. As a result, crafted perturbations may be dispersed onto irrelevant background regions, diluting both stealth and cross-domain/model efficacy.
@@ 2.1 Transfer-based Adversarial Attacks
Transfer-based attacks exploit the empirical finding that adversarial perturbations crafted on one model often remain effective against others, even when architectures or training data differ. Early methods relied on iterative gradient-based strategies—momentum-integrated attacks (DI [77], TI [16]), input-diversity techniques (SI [39], Admix [71]), and strong baselines such as BIM [34], PGD [43], and C&W [9]. These approaches enhance transferability via gradient smoothing, input transformations, and ensemble gradients, but incur heavy per-example optimization costs and often struggle against architectures that diverge significantly from the surrogate.
More recent work has introduced generative frameworks that train feed-forward generators to synthesize perturbations in a single pass. GAP [51], CDA [50], and LTP [29] demonstrated orders-of-magnitude speedups with comparable transfer rates. Subsequent advances—BIA [86], GAMA [1], FACL-Attack [81], PDCL-Attack [83], and NAT [48]—have further improved robustness by integrating logit or mid-level layer feature divergence, frequency-domain constraints, text prompt-driven, and neuron-targeted losses. Our distinctions. Rather than focusing solely on end-to-end optimization or domain-level constraints, we analyze the generator’s intermediate feature hierarchy and preserve semantic fidelity in its early blocks to steer perturbations onto object-centric regions, thereby enhancing cross-model transfer effectiveness.
@@ 2.2 Generative Model-based Attacks
Generative attacks recast adversarial synthesis as a learning problem, training an image-to-image network (e.g. GAN or encoder–decoder) to produce perturbations in one pass. GAP [51] pioneered a framework in which the generator outputs adversarial noise that is then added to the input. CDA [70] extends this by training a transformation network that directly outputs adversarial examples. Subsequent works incorporate perceptual losses based on surrogate logits [29] and mid-level surrogate features [86, 48]. Building on the feature-similarity loss of Zhang et al. [86], more recent approaches leverage vision–language foundation models such as CLIP [52] [1, 83] and apply frequency-domain manipulations to surrogate features [82], further boosting transferability. Our distinctions. While prior frameworks prioritize perturbation realism or frequency characteristics, we explicitly target the generator’s internal semantics by combining Mean Teacher-based smoothing with feature distillation on early blocks, preserving object contours and textures and concentrating adversarial perturbations in the most transferable regions.
@@ 2.3 Self-Knowledge Distillation
Self-knowledge distillation (Self-KD) aims to train a model to refine its own representations without an external teacher. Pioneering works in this field, Born-Again Networks [19] and Deep Mutual Learning [87], demonstrated that iterative self- and peer-distillation can improve generalization and robustness. Recent works [38, 84] incorporate self-kd by aligning logits or intermediate features within its own network, or progressively updating the network [30]. In this paradigm of using a student-teacher framework, the Mean Teacher framework [62], originally developed for semi-supervised learning, aims to maintain a teacher as the exponential moving average of the student’s weights, implicitly enforcing temporal consistency in predictions or feature maps. This EMA-based smoothing has been shown to reduce overfitting, stabilize training, and enhance domain invariance—properties that are directly relevant to generating perturbations that transfer across black-box models. Our distinctions. Departing from classification-centric distillation, we integrate the Mean Teacher paradigm into a generative attack pipeline, using EMA to smooth intermediate features and applying hinge-based feature distillation on early blocks to preserve semantic integrity critical for cross-setting transferability.
@@ 3.1 Problem Formulation
Our framework for generative model-based untargeted transfer attacks comprises an adversarial perturbation generator $\mathcal{G}_{\theta}(\cdot)$, producing unconstrained adversarial examples $x^{adv}$ from benign inputs $x$, which are then projected via a perturbation projector $\mathcal{P}(\cdot)$ to satisfy $||\mathcal{P}(x^{adv})-x||_{\infty}\leq\epsilon$. To train $\mathcal{G}_{\theta}(\cdot)$ in a label-free manner in the untargeted attack, we employ a white-box surrogate model to provide an adversarial supervisory signal for generator updates via back-propagation. The adversarial loss uses mid-layer surrogate features $\mathcal{F}_{k}(\cdot)$, which contain model-shared characteristics that are crucial for adversarial transferability.
@@ 3.2 Motivation
Our motivation draws from an empirical observation of how perturbations evolve through the residual blocks within the generator after training (see Fig. 3). Given the ResNet-based generator architecture by previous works [50, 29, 86, 1, 82, 83], we visualized the output activation maps from the intermediate layers, as most of the noise is added in the residual intermediate blocks as pointed out by [86].
@@ 3       Randomly sample a mini-batch $x_{i}$ from $\mathcal{D}_{\textrm{train}}$
7       Project $x_{i}^{\textrm{adv}}$ using the projector $\mathcal{P}$ within the perturbation budget: $||\mathcal{P}(x_{i}^{\textrm{adv}})-x_{i}||_{\infty}\leq\epsilon$
@@ 3.3 Proposed Framework
Our semantic structure–aware attack, as described in Alg. 1, augments a standard generative adversarial attack with two key components: a Mean Teacher-based feature smoothing and a self-feature distillation loss that enforces semantic consistency in early generator layers. We base our approach on the baseline work [86] because all the subsequent works [1, 82, 83] base their losses on the feature similarity-based adversarial loss suggested by [86], thus it is adequate to serve as a solid baseline.
@@
The Mean Teacher (MT) framework [62, 13, 37, 88, 7, 14] has consistently demonstrated robustness in tasks characterized by significant domain shifts between training and testing. Its core mechanism of updating the teacher’s parameters with EMA of the student’s parameters provides a form of temporal ensembling that naturally suppresses instance-specific noise. Intuitively, this EMA update smooths out high-frequency perturbation artifacts, enriching the semantic consistency and stability of the teacher’s intermediate feature maps. As a result, these smoothed features serve as a reliable reference for the student, helping to preserve object contours and shapes throughout adversarial synthesis. To integrate MT, we maintain two generators: a student $G_{\theta}$ that is trained via gradient descent, and a teacher $G_{\theta^{\prime}}$ whose parameters are updated by Eq. 1. We set these mean teacher features as a reference for our distillation component.
@@
Object-salient intermediate representations have been shown to be critical for adversarial transfer in black-box settings [73, 5, 31, 86], and recent work has explored manipulating input or surrogate-level features to this end [27, 36, 48]. In our generative framework, however, a naïve generator progressively loses semantic integrity in its intermediate layers (Fig. 2), scattering perturbations away from object-salient regions. To preserve these crucial object cues, we introduce a self-distillation mechanism grounded in the Mean Teacher paradigm [62, 21, 10, 35].
@@
where $\texttt{cos}(\cdot,\cdot)$ is cosine similarity and $\tau$ a similarity threshold. This loss anchors the student’s edges and shape priors to the teacher’s smoothed semantics, ensuring that subsequent perturbations concentrate on object-centric regions. $\mathcal{W}_{\textrm{distill}}\in\mathbb{R}^{|L|}$ denotes the softmax output of a learnable parameter for intermediate block-wise loss weighting. When combined with the adversarial objective, this self-feature distillation yields perturbations that are both highly transferable and tightly aligned with the image’s core structure. For fair comparisons with the state-of-the-art methods, we adopt the adversarial loss in the surrogate feature space as practiced in the baseline [86].
@@ 4.1 Implementation details
Throughout the experiments, we train the perturbation generator with $\epsilon=10$ using data from ImageNet-1K [55] containing 1.2 M natural images of 224$\times$224$\times$3 resolution, following [51, 50, 29, 86, 1, 81, 82, 83]. We also use the mid-level layer feature at $k=16$ (Maxpooling.3) of VGG-16 surrogate for our baseline [86]. We set $\lambda_{distill}=0.7$ throughout our experiments for stable generator training at the feature level, and the EMA update parameter $\eta=0.999$ following [62]. We selected $L_{early}=\{1,2\}$ for distilling the first and second intermediate residual blocks within the generator. We compare our attacks against the state-of-the-art baselines that rely on the same ResNet-based generator to craft adversarial examples, i.e., CDA [50], LTP [29], BIA [86], GAMA [1], FACL-Attack [81, 82], and PDCL-Attack [83].
@@ 4.2 Evaluation settings
We evaluate adversarial transfer under two black-box protocols. In the cross-model setting, perturbations are crafted on surrogate models trained with the same data distribution (ImageNet-1K [55]) and then tested on unseen target architectures. In the cross-domain/task settings, adversarial examples are to generalize across domain/task shifts without access to any target-distribution samples.
@@
For cross-model evaluation, we employ ImageNet-1K ($224\times 224$ resolution, 1,000 classes) pre-trained classification models of various architectures with their publicly available model weights. We source the pre-trained models from TorchVision [46] and Timm [72] libraries. Compared to previous approaches [29, 82, 83] demonstrating cross-model architecture transferability, we expand the evaluation to a wider scope of target model architectures for enhanced architecture-agnostic transferability. We test our attack against a total of 22 different model architectures (12 CNN-based [24, 26, 59, 53, 61, 28, 60, 69], 6 ViT-based [66, 41, 64, 4, 6], 2 Mixers [63, 65], and 2 Vision Mamba-based [89, 23]) for cross-model evaluations as in Table 1.
For cross-domain evaluation, we validate our attack against three different models (i.e., ResNet50 [24], SE-Net and SE-ResNet101 [25]) pre-trained on fine-grained datasets, CUB-200-2011 [67] (200 classes), Stanford Cars [33] (196 classes), FGVC Aircraft [45] (100 classes), of $448\times 448$ resolution. We detail statistics and descriptions in the Appendix. For cross-task, we select a CNN-based and a ViT-based model for each task of semantic segmentation (SS) and object detection (OD), whose pre-trained weights are openly accessible as with the ImageNet-1K pre-trained weights. Specifically, we test against DeepLabV3+ [11] and SegFormer [78] for SS and Faster R-CNN [20] on detectron2 and DETR [8] for OD. We validate on Cityscapes [12] and COCO’17 [40] for SS and OD tasks, respectively.
@@
We also test our attack against robust models, i.e., adversarially trained Inception-V3 [34], ViT [18] and ConvNeXt [57] models, and robust input processing methods such as JPEG (75%) [22], bit reduction (BDR; 4-bit) [80] and randomization (R&P) [75] in Table 8.
Table 1: Quantitative cross-model transferability results. We report the average improvement margins of our method added to each baseline, averaged over three models for each domain. We report the improvements ($\Delta$p) with ours relative to each baseline. Better averaged results are marked in boldface.
@@
BIA [86] w/ Ours Acc. ($\Delta$%p) $\downarrow$ -2.23 -1.99 -1.29 0.01 -3.72 -1.59 -3.28 -2.85 -0.40 -2.53 -2.70 -0.94 -0.67 -0.25 -11.99 -1.34 -0.40 -1.08 -1.50 -4.32 -2.53 0.00 -2.16
@@
Table 2: Quantitative cross-domain transferability results. We report the average improvement ($\Delta$p) with ours added from each baseline for each domain, in gray row. Better results in boldface.
Cross-domain CUB-200-2011 [67] Stanford Cars [33] FGVC Aircraft [45] Avg. Acc.
@@
Table 3: Quantitative cross-task transferability results. We report the average improvement ($\Delta$p) for our components applied to different generator architectures and evaluated against semantic segmentation (mIoU $\downarrow$) and object detection (mAP50 $\downarrow$) models. MT denotes mean teacher, and better results in boldface.
@@
Method Benign Baseline +MT +MT+$\mathcal{L}_{\textrm{distill}}$ Baseline +MT +MT+$\mathcal{L}_{\textrm{distill}}$
SS DeepLabV3+ [11] 76.21 24.22 +0.76 -2.92 23.89 -0.79 -1.84
SegFormer [78] 71.89 29.34 -3.31 -3.81 25.60 -0.78 -0.85
@@
OD FRCNN [20] 61.01 27.51 -0.05 -0.12 28.43 +0.03 -0.09
DETR [8] 62.36 23.92 -3.18 -3.54 21.01 -0.02 -0.29
@@
While developing an effective attack mechanism is crucial, it must be validated by fair and comprehensive evaluations. Current evaluation protocols adopted by prior works [51, 50, 29, 86, 1, 82, 83, 48] exhibit three key limitations. First, most studies report only a single primary metric (either one of Attack Success Rate (ASR), Fooling Rate (FR), or Top-1 Accuracy (%)), offering only a one-dimensional view of attack robustness and neglecting other aspects such as unintended corrections in predictions. Second, datasets and sample sizes are often selected arbitrarily or limited to a single scale, preventing a fair comparison across attacks and undermining statistical significance. Third, evaluations in prior works commonly target a narrow set of victim architectures (e.g., mostly CNN-based), lacking the diversity of modern model families, including vision transformers (ViTs) and state-space models (SSMs), and thus overstating robustness. To address these shortcomings, we introduce the Accidental Correction Rate (ACR) as a complementary metric that captures the proportion of adversarial inputs that inadvertently restore correct predictions, enriching the assessment of attack efficacy alongside conventional measures (i.e., ASR, FR, Accuracy).
We evaluate the effectiveness and transferability of adversarial attacks across model architectures and domain shifts using four key metrics. For notational convenience, let $f(x)$ denote the predicted label for input $x$, $f(x+\delta)$ the prediction after applying adversarial perturbation $\delta$, and $y$ the ground-truth label. The evaluation set is denoted by $\mathcal{D}$, with $\mathcal{C}=\{x\in\mathcal{D}\mid f(x)=y\}$ representing correctly classified samples, and =$\mathcal{I}=\{x\in\mathcal{D}\mid f(x)\neq y\}$ denoting misclassified samples under clean inference. We formally define our evaluation metrics as follows:

### v4

@@ 1 Introduction
Deep neural networks have driven transformative advances across computer vision, natural language processing, and medical diagnosis. Their capacity to learn rich, hierarchical feature representations has enabled breakthroughs in tasks ranging from fine-grained image classification to complex sequence generation. Yet these models remain alarmingly fragile: even small, human-imperceptible perturbations—known as adversarial examples (AEs) Szegedy et al. (2013)—can provoke confident misclassification, underscoring the need to safeguard real-world deployments against malicious exploits. The threat is further intensified in black-box scenarios, where attackers have no access to a model’s architecture or parameters. In such scenarios, transferable adversarial attacks craft perturbations on a surrogate network and deploy the same perturbation against unknown targets. This capability empowers adversaries to compromise multiple real-world systems, such as autonomous vehicles, biometric scanners, and diagnostic tools, with a single perturbation strategy, posing severe risks to safety-critical applications.
@@
Early white-box iterative attacks, such as FGSM Zhang et al. (2021) and its iterative variants Dong et al. (2018); Xie et al. (2019a); Dong et al. (2019b), exploit gradient information directly but inherently require full model access. To evaluate real-world attack robustness, transfer-based (black-box) attacks emerged, leveraging the phenomenon that adversarial perturbations crafted on one model often fool others. While iterative gradient-based methods Madry et al. (2017a); Carlini et al. (2019); Zhang et al. (2021), whether operating in white-box or query-based modes, can yield moderately transferable adversarial examples, they incur intensive costs due to per-sample optimization steps, resulting in prohibitive computation time and slow inference. Such bottleneck prohibits their use in time-sensitive scenarios and limits scalable and practically deployable attacks.
@@
To overcome this, generative transfer attacks Xiao et al. (2018); Wang et al. (2018); Baluja and Fischer (2017, 2018); Poursaeed et al. (2018); Naseer et al. (2019); kanth Nakka and Salzmann (2021); Zhang et al. (2022); Aich et al. (2022); Yang et al. (2023, 2024a, 2024b); Nakka and Alahi (2025) aim to train a dedicated perturbation generator against a surrogate, enabling fast inference and high transfer rates without per-example optimization. Generative attacks address these limitations by training a feed-forward generator to produce perturbations in one pass. Once trained, this generator yields adversarial noise with a single forward inference, offering orders-of-magnitude improvements in speed, seamless scalability to large datasets, and strong transferability. However, existing generative frameworks predominantly treat the generator as a black box, optimizing only end-to-end metrics while overlooking how semantic information, such as object edges and coarse shapes, is represented and potentially degraded within its intermediate layers. As a result, crafted perturbations may be dispersed onto irrelevant background regions, diluting both stealth and cross-domain/model efficacy.
@@ 2.1 Transfer-based Adversarial Attacks
Transfer-based attacks exploit the empirical finding that adversarial perturbations crafted on one model often remain effective against others, even when architectures or training data differ. Early methods relied on iterative gradient-based strategies—momentum-integrated attacks (DI Xie et al. (2019b), TI Dong et al. (2019a)), input-diversity techniques (SI Lin et al. (2019), Admix Wang et al. (2021)), and strong baselines such as BIM Kurakin et al. (2016), PGD Madry et al. (2017a), and C&W Carlini et al. (2019). These approaches enhance transferability via gradient smoothing, input transformations, and ensemble gradients, but incur heavy per-example optimization costs and often struggle against architectures that diverge significantly from the surrogate.
More recent work has introduced generative frameworks that train feed-forward generators to synthesize perturbations in a single pass. GAP Poursaeed et al. (2018), CDA Naseer et al. (2019), and LTP kanth Nakka and Salzmann (2021) demonstrated orders-of-magnitude speedups with comparable transfer rates. Subsequent advances—BIA Zhang et al. (2022), GAMA Aich et al. (2022), FACL-Attack Yang et al. (2023), PDCL-Attack Yang et al. (2024b), and NAT Nakka and Alahi (2025)—have further improved robustness by integrating logit or mid-level layer feature divergence, frequency-domain constraints, text prompt-driven, and neuron-targeted losses. Our distinctions. Rather than focusing solely on end-to-end optimization or domain-level constraints, we analyze the generator’s intermediate feature hierarchy and preserve semantic fidelity in its early blocks to steer perturbations onto object-centric regions, thereby enhancing cross-model transfer effectiveness.
@@ 2.2 Generative Model-based Attacks
Generative attacks recast adversarial synthesis as a learning problem, training an image-to-image network (e.g. GAN or encoder–decoder) to produce perturbations in one pass. GAP Poursaeed et al. (2018) pioneered a framework in which the generator outputs adversarial noise that is then added to the input. CDA Wang et al. (2022) extends this by training a transformation network that directly outputs adversarial examples. Subsequent works incorporate perceptual losses based on surrogate logits kanth Nakka and Salzmann (2021) and mid-level surrogate features Zhang et al. (2022); Nakka and Alahi (2025). Building on the feature-similarity loss of Zhang et al. Zhang et al. (2022), more recent approaches leverage vision–language foundation models such as CLIP Radford et al. (2021) Aich et al. (2022); Yang et al. (2024b) and apply frequency-domain manipulations to surrogate features Yang et al. (2024a), further boosting transferability. Our distinctions. While prior frameworks prioritize perturbation realism or frequency characteristics, we explicitly target the generator’s internal semantics by combining Mean Teacher-based smoothing with feature distillation on early blocks, preserving object contours and textures and concentrating adversarial perturbations in the most transferable regions.
@@ 2.3 Self-Knowledge Distillation
Self-knowledge distillation (Self-KD) aims to train a model to refine its own representations without an external teacher. Pioneering works in this field, Born-Again Networks Furlanello et al. (2018) and Deep Mutual Learning Zhang et al. (2018), demonstrated that iterative self- and peer-distillation can improve generalization and robustness. Recent works Li et al. (2024); Yun et al. (2020) incorporate self-kd by aligning logits or intermediate features within its own network, or progressively updating the network Kim et al. (2021). In this paradigm of using a student-teacher framework, the Mean Teacher framework Tarvainen and Valpola (2017), originally developed for semi-supervised learning, aims to maintain a teacher as the exponential moving average of the student’s weights, implicitly enforcing temporal consistency in predictions or feature maps. This EMA-based smoothing has been shown to reduce overfitting, stabilize training, and enhance domain invariance—properties that are directly relevant to generating perturbations that transfer across black-box models. Our distinctions. Departing from classification-centric distillation, we integrate the Mean Teacher paradigm into a generative attack pipeline, using EMA to smooth intermediate features and applying hinge-based feature distillation on early blocks to preserve semantic integrity critical for cross-setting transferability.
@@ 3.1 Problem Formulation
Our framework for generative model-based untargeted transfer attacks comprises an adversarial perturbation generator $\mathcal{G}_{\theta}(\cdot)$, producing unconstrained adversarial examples $x^{adv}$ from benign inputs $x$, which are then projected via a perturbation projector $\mathcal{P}(\cdot)$ to satisfy $\mathcal{P}(x^{adv})-x||_{\infty}\leq\epsilon$. To train $\mathcal{G}_{\theta}(\cdot)$ in a label-free manner in the untargeted attack, we employ a white-box surrogate model to provide an adversarial supervisory signal for generator updates via back-propagation. The adversarial loss uses mid-layer surrogate features $\mathcal{F}_{k}(\cdot)$, which contain model-shared characteristics that are crucial for adversarial transferability.
@@ 3.2 Motivation
Our motivation draws from an empirical observation of how perturbations evolve through the residual blocks within the generator after training (see Fig. 3). Given the ResNet-based generator architecture by previous works Naseer et al. (2019); kanth Nakka and Salzmann (2021); Zhang et al. (2022); Aich et al. (2022); Yang et al. (2024a, b), we visualized the output activation maps from the intermediate layers, as most of the noise is added in the residual intermediate blocks as pointed out by Zhang et al. (2022).
@@ 3   Randomly sample a mini-batch $x_{i}$ from $\mathcal{D}_{\textrm{train}}$
7   Project $x_{i}^{\textrm{adv}}$ using the projector $\mathcal{P}$ within the perturbation budget: $\mathcal{P}(x_{i}^{\textrm{adv}})-x_{i}||_{\infty}\leq\epsilon$
@@ 3.3 Proposed Framework
Our semantic structure–aware attack, as described in Alg. 1, augments a standard generative adversarial attack with two key components: a Mean Teacher-based feature smoothing and a self-feature distillation loss that enforces semantic consistency in early generator layers. We base our approach on the baseline work Zhang et al. (2022) because all the subsequent works Aich et al. (2022); Yang et al. (2024a, b) base their losses on the feature similarity-based adversarial loss suggested by Zhang et al. (2022), thus it is adequate to serve as a solid baseline.
@@
The Mean Teacher (MT) framework Tarvainen and Valpola (2017); Deng et al. (2021); Li et al. (2022); Zhao et al. (2022); Cao et al. (2023); Döbler et al. (2023) has consistently demonstrated robustness in tasks characterized by significant domain shifts between training and testing. Its core mechanism of updating the teacher’s parameters with EMA of the student’s parameters provides a form of temporal ensembling that naturally suppresses instance-specific noise. Intuitively, this EMA update smooths out high-frequency perturbation artifacts, enriching the semantic consistency and stability of the teacher’s intermediate feature maps. As a result, these smoothed features serve as a reliable reference for the student, helping to preserve object contours and shapes throughout adversarial synthesis. To integrate MT, we maintain two generators: a student $G_{\theta}$ that is trained via gradient descent, and a teacher $G_{\theta^{\prime}}$ whose parameters are updated by Eq. 1. We set these mean teacher features as a reference for our distillation component.
@@
Object-salient intermediate representations have been shown to be critical for adversarial transfer in black-box settings Wu et al. (2020); Byun et al. (2022); Kim et al. (2022); Zhang et al. (2022), and recent work has explored manipulating input or surrogate-level features to this end Huang et al. (2019); Li et al. (2023); Nakka and Alahi (2025). In our generative framework, however, a naïve generator progressively loses semantic integrity in its intermediate layers (Fig. 2), scattering perturbations away from object-salient regions. To preserve these crucial object cues, we introduce a self-distillation mechanism grounded in the Mean Teacher paradigm Tarvainen and Valpola (2017); Grill et al. (2020); Caron et al. (2021); Lee et al. (2023).
@@
where $\texttt{cos}(\cdot,\cdot)$ is cosine similarity and $\tau$ a similarity threshold. This loss anchors the student’s edges and shape priors to the teacher’s smoothed semantics, ensuring that subsequent perturbations concentrate on object-centric regions. $\mathcal{W}_{\textrm{distill}}\in\mathbb{R}^{|L|}$ denotes the softmax output of a learnable parameter for intermediate block-wise loss weighting. When combined with the adversarial objective, this self-feature distillation yields perturbations that are both highly transferable and tightly aligned with the image’s core structure. For fair comparisons with the state-of-the-art methods, we adopt the adversarial loss in the surrogate feature space as practiced in the baseline Zhang et al. (2022).
@@ 4.1 Implementation details
Throughout the experiments, we train the perturbation generator with $\epsilon=10$ using data from ImageNet-1K Russakovsky et al. (2015) containing 1.2 M natural images of $\times224$$\times$3 resolution, following Poursaeed et al. (2018); Naseer et al. (2019); kanth Nakka and Salzmann (2021); Zhang et al. (2022); Aich et al. (2022); Yang et al. (2023, 2024a, 2024b). We also use the mid-level layer feature at $k=16$ (Maxpooling.3) of VGG-16 surrogate for our baseline Zhang et al. (2022). We set $\lambda_{distill}=0.7$ throughout our experiments for stable generator training at the feature level, and the EMA update parameter $\eta=0.999$ following Tarvainen and Valpola (2017). We selected $L_{early}=\{1,2\}$ for distilling the first and second intermediate residual blocks within the generator. We compare our attacks against the state-of-the-art baselines that rely on the same ResNet-based generator to craft adversarial examples, i.e., CDA Naseer et al. (2019), LTP kanth Nakka and Salzmann (2021), BIA Zhang et al. (2022), GAMA Aich et al. (2022), FACL-Attack Yang et al. (2023, 2024a), and PDCL-Attack Yang et al. (2024b).
@@ 4.2 Evaluation settings
We evaluate adversarial transfer under two black-box protocols. In the cross-model setting, perturbations are crafted on surrogate models trained with the same data distribution (ImageNet-1K Russakovsky et al. (2015)) and then tested on unseen target architectures. In the cross-domain/task settings, adversarial examples are to generalize across domain/task shifts without access to any target-distribution samples.
@@
For cross-model evaluation, we employ ImageNet-1K ($224\times 224$ resolution, 1,000 classes) pre-trained classification models of various architectures with their publicly available model weights. We source the pre-trained models from TorchVision Marcel and Rodriguez (2010) and Timm Wightman libraries. Compared to previous approaches kanth Nakka and Salzmann (2021); Yang et al. (2024a, b) demonstrating cross-model architecture transferability, we expand the evaluation to a wider scope of target model architectures for enhanced architecture-agnostic transferability. We test our attack against a total of 22 different model architectures (12 CNN-based He et al. (2016); Huang et al. (2017); Szegedy et al. (2016); Radosavovic et al. (2020); Tan et al. (2019); Iandola et al. (2016); Tan and Le ; Wang et al. (2020), 6 ViT-based Tu et al. (2022); Liu et al. (2021); Touvron et al. (2021); Bao et al. (2021); Cai et al. (2022), 2 Mixers Tolstikhin et al. (2021); Trockman and Kolter (2022), and 2 Vision Mamba-based Zhu et al. (2024); Hatamizadeh and Kautz (2025)) for cross-model evaluations as in Table 1.
For cross-domain evaluation, we validate our attack against three different models (i.e., ResNet50 He et al. (2016), SE-Net and SE-ResNet101 Hu et al. (2018)) pre-trained on fine-grained datasets, CUB-200-2011 Wah et al. (2011) (200 classes), Stanford Cars Krause et al. (2013) (196 classes), FGVC Aircraft Maji et al. (2013) (100 classes), of $448\times 448$ resolution. We detail statistics and descriptions in the Appendix. For cross-task, we select a CNN-based and a ViT-based model for each task of semantic segmentation (SS) and object detection (OD), whose pre-trained weights are openly accessible as with the ImageNet-1K pre-trained weights. Specifically, we test against DeepLabV3+ Chen et al. (2018) and SegFormer Xie et al. (2021) for SS and Faster R-CNN Girshick (2015) on detectron2 and DETR Carion et al. (2020) for OD. We validate on Cityscapes Cordts et al. (2016) and COCO’17 Lin et al. (2014) for SS and OD tasks, respectively.
@@
We also test our attack against robust models, i.e., adversarially trained Inception-V3 Kurakin et al. (2016), ViT Dosovitskiy et al. (2021) and ConvNeXt Singh et al. (2023) models, and robust input processing methods such as JPEG (75%) Guo et al. (2017), bit reduction (BDR; 4-bit) Xu et al. (2018) and randomization (R&P) Xie et al. (2018) in Table 8.
Table 1: Quantitative cross-model transferability results. We report the average improvement margins of our method added to each baseline, averaged over three models for each domain. We report the improvements ($\Delta$ %p) with ours relative to each baseline. Better averaged results are marked in boldface.
@@
BIA Zhang et al. (2022) w/ Ours Acc. ($\Delta$%p) $\downarrow$ -2.23 -1.99 -1.29 0.01 -3.72 -1.59 -3.28 -2.85 -0.40 -2.53 -2.70 -0.94 -0.67 -0.25 -11.99 -1.34 -0.40 -1.08 -1.50 -4.32 -2.53 0.00 -2.16
@@
Table 2: Quantitative cross-domain transferability results. We report the average improvement ($\Delta$ %p) with ours added from each baseline for each domain, in gray row. Better results in boldface.
Cross-domain CUB-200-2011 Wah et al. (2011) Stanford Cars Krause et al. (2013) FGVC Aircraft Maji et al. (2013) Avg. Acc.
@@
Table 3: Quantitative cross-task transferability results. We report the average improvement ($\Delta$ %p) for our components applied to different generator architectures and evaluated against semantic segmentation (mIoU $\downarrow$) and object detection (mAP50 $\downarrow$) models. MT denotes mean teacher, and better results in boldface.
@@
Method Benign $\mathcal{L}_{\textrm{distill}} Baseline +MT +MT$$\mathcal{L}_{\textrm{distill}}$
$S$
DeepLabV3+ Chen et al. (2018) 76.21 24.22 +0.76 -2.92 23.89 -0.79 -1.84
SegFormer Xie et al. (2021) 71.89 29.34 -3.31 -3.81 25.60 -0.78 -0.85
@@
OD
FRCNN Girshick (2015) 61.01 27.51 -0.05 -0.12 28.43 +0.03 -0.09
DETR Carion et al. (2020) 62.36 23.92 -3.18 -3.54 21.01 -0.02 -0.29
@@
While developing an effective attack mechanism is crucial, it must be validated by fair and comprehensive evaluations. Current evaluation protocols adopted by prior works Poursaeed et al. (2018); Naseer et al. (2019); kanth Nakka and Salzmann (2021); Zhang et al. (2022); Aich et al. (2022); Yang et al. (2024a, b); Nakka and Alahi (2025) exhibit three key limitations. First, most studies report only a single primary metric (either one of Attack Success Rate (ASR), Fooling Rate (FR), or Top-1 Accuracy (%)), offering only a one-dimensional view of attack robustness and neglecting other aspects such as unintended corrections in predictions. Second, datasets and sample sizes are often selected arbitrarily or limited to a single scale, preventing a fair comparison across attacks and undermining statistical significance. Third, evaluations in prior works commonly target a narrow set of victim architectures (e.g., mostly CNN-based), lacking the diversity of modern model families, including vision transformers (ViTs) and state-space models (SSMs), and thus overstating robustness. To address these shortcomings, we introduce the Accidental Correction Rate (ACR) as a complementary metric that captures the proportion of adversarial inputs that inadvertently restore correct predictions, enriching the assessment of attack efficacy alongside conventional measures (i.e., ASR, FR, Accuracy).
We evaluate the effectiveness and transferability of adversarial attacks across model architectures and domain shifts using four key metrics. For notational convenience, let $f(x)$ denote the predicted label for input $x$, $f(x+\delta)$ the prediction after applying adversarial perturbation $\delta$, and $y$ the ground-truth label. The evaluation set is denoted by $\mathcal{D}$, with $\mathcal{C}=\{x\in\mathcal{D}\mid f(x)=y\}$ representing correctly classified samples, and $\mathcal{I}=\{x\in\mathcal{D}\mid f(x)\neq y\}$ denoting misclassified samples under clean inference. We formally define our evaluation metrics as follows: